# Multi–charged geometries with cosmological constant

### v1

@@ 3.1 Solution for a single stack of multi–charged $p$–branes
$S=\int dx\sqrt{-g}\left[R+\frac{N_{p}\gamma^{2}\nu}{4[n_{-}\nu-2]}(\partial\psi)^{2}-\frac{e^{\gamma\psi}}{2}F_{(p+2)}\wedge\star F_{(p+2)}\right]\,.$ (3.7)
@@
    Assuming that $n_{-}\nu\neq 2$, one can rescale the scalar field to choose a convenient normalization: There are two natural options: the canonical normalization that would give
@@
    We will choose the latter option, which gives the kinetic term $\frac{4}{d-2}(\partial\psi)^{2}$. Then comparing (3.7) with the action of type IIB supergravity [46], one can embed strings and various branes:
@@
    Since the geometry (3.1) has multiple gauge fields, it covers not only only individual branes, but some smeared brane intersections as well. Specifically, let us consider a configuration D(p+4) branes stretched in the directions $(t,x^{1},\dots,x^{p+4})$ and Dp branes stretched in $(t,x^{1},\dots,x^{p})$. In the string frame, the metric reads
@@
    Upon smearing the lower dimensional branes in $(x^{p+1},\dots,x^{p+4})$ directions and reducing the system along this four–dimensional torus, one ends up with an effective supergravity action (3.1) with two gauge fields: one is sourced by D(p+4) branes, and another one is sourced by the Dp branes. Generically, there are also two scalars: the original dilaton and the volume factor of the 4–torus. Dimensional reduction from type II supergravity gives
@@
    $F_{5}$ $=$ $\frac{1}{4}dt\wedge d\left[\frac{1}{h_{1}}dX_{123}+\frac{1}{h_{2}}dX_{1}dY_{23}-\frac{1}{h_{3}}dX_{2}dY_{13}+\frac{1}{h_{4}}dX_{3}dY_{12}\right]+dual,$ (3.21)
    $h_{i}$ $=$ $1+\frac{Q_{i}}{r},\quad f=h_{1}h_{2}h_{3}h_{4}\,.$
@@
    we conclude the scalar are subject to constraint $\sum\psi_{i}=0$ due to the relation $k\nu n_{-}=2$. As we will show in the next subsection, this has important implications for the near–horizon limit of the geometry (f).
@@ 3.2 Branes in the near–horizon limit and wormhole geometries
In the last subsection we constructed the geometries (3.1) which describe multi–charged black branes with regular horizons located at $r_{h}=2m$. As in the Schwarzschild case, one encounters a coordinate singularity at $r=r_{h}$, and the space can be continued using Kruskal coordinates. In this subsection we will focus on the extremal case of the geometries (3.1), for which the horizon is pushed to $r=0$, and explore their near–horizon limits. We will see that for a particular choice of the parameter $\nu$, such near horizon limit gives a product of the $n$–dimensional sphere and an AdS space on the Poincare patch according to the same mechanism that led to the discovery of the AdS/CFT correspondence [5]. Going to the global coordinates on AdS, one removes the coordinate singularity at the origin, but also cuts out the asymptotically–flat region [6]. However, such a transition also allows one to introduce a cosmological constant or a nontrivial potential for scalars and extend the solutions (3.1) to geometries with AdSp+n+2 rather than flat asymptotics. In this subsection we will explore such modifications, generalizing the singly-charged wormhole–type geometries which were constructed in [19].
@@
The fluxes are still given by (3.1). The solutions (3.2) have curvature singularities at $r=0$ unless $k\nu n_{-}=2$: the easiest way to see this is to observe the the scalars make divergent contributions to the stress–energy tensor, but singularities can also be found by direct calculations. The regular solutions with $k\nu n_{-}=2$ have the following properties:
@@
    which holds everywhere, not only in the near horizon limit. Therefore, we conclude that to avoid curvature singularities in solutions, the action (3.1) must have $k$ gauge fields but only $k-1$ independent scalars.
@@
    Some string theory examples of geometries with the regularity condition $k\nu n_{-}=2$ were reviewed in (3.1), and not surprisingly, they contain the extremal D3, M2, M5 branes, D1–D5 system, and D3–D3–D3–D3 intersections. It would be interesting to find physical applications of other regular cases as well. Note that many string theory solutions that don’t satisfy the regularity condition $k\nu n_{-}=2$, such as geometries produced by fundamental strings and D branes, are allowed in the full theory and lead to singularities only in the supergravity approximation.
@@
We begin with reviewing the results for the $k=1$ case, which were obtained in [19]. In this case, there is only one gauge field and no scalars due to the constant (3.27) so the potential is representing the constant everywhere. Both the geometry (3.29) and the asymptotic AdSp+n+2 have the isometries of an $n$–dimensional sphere and a $p+1$–dimensional AdS space, so we will assume that the connecting solution has these isometries as well. This leads to an ansatz for the metric and the field strength:
@@
As demonstrated above, regular geometries can be obtained only if the condition $k\nu n_{-}=2$ is satisfied, and this imposes the constraint (3.27) on scalar fields. For such configurations, the the ansatz (3.31) has a unique extension consistent with symmetries:
@@
Index $I$ runs from $1$ to $(k-1)$, and the values of $e^{(I)}_{0}$ cab be extracted from (3.26). These $(k+1)$ expansions can be substituted into three Einstein’s equations and $(k-1)$ equations for the scalars, and as in the case of one charge (3.31), one of Einstein’s equations turns out to be redundant. Therefore, at each order in $r$, one finds $(k+1)$ linear equations for $(k+1)$ variables, leading to the unique solution. The only exception from this rule is zeroth order, when one encounters a nonlinear constraint on $(A_{0},C_{0})$, which is solved by a straightforward counterpart of (3.33):
@@
In particular, one concludes that solutions (3.35) regular at $r=0$ are uniquely specified by one parameter $C_{0}$. Beyond the perturbative expansions (3.36) one can perform numerical integration of the system and analyze the behavior of solutions at infinity. One finds a phase structure which is qualitatively similar to items (c)–(f) above with one major difference: now the scalar fields $\psi_{I}$ have nontrivial profiles. We conclude that an asymptotic behavior from each of the options (d)-(f) above lead to a unique regular solution (3.35). The numerical values $(C_{\star},C_{\bullet})$ depend on the number of scalar fields and values of charges.
@@ 3.3 Smeared branes and brane intersections
In section 3.1 we constructed geometries (3.1) which solve equations of motion coming from the action (3.1). To do so, we imposed the ansatz (3.1) in which the number of flat directions was correlated with the rank of the field strength. Specifically, all flat directions were longitudinal to the branes. In this subsection we will extend the earlier results in two directions: we will construct the geometries describing smeared branes, in which some of the flat directions are longitudinal to the worldvolume, and the solutions produced by branes within branes. Some examples of second types of solutions were briefly discussed in section 3.1, where we analyzed their reductions to lower dimensions, and here we will construct the full uplifted geometries. In contrast to the string theory examples discussed in section 3.1, generically we will be able to embed $s$–branes inside $p$ branes for any values of $(s,p)$, and the well known relation $p=s+4$ is recovered when the coupling of scalars are adjusted to their string theoretic values.
@@
$S$ $=$ $\int d^{d}x\sqrt{-g}\left[R-\sum_{I}^{k}\frac{e^{\gamma\psi_{I}}}{2}F_{(p+2)}^{I}\wedge\star F_{(p+2)}^{I}-\sum_{J}^{\bar{k}}\frac{e^{\bar{\gamma}{\bar{\psi}}_{J}}}{2}{\bar{F}}_{(s+2)}^{J}\wedge\star{\bar{F}}_{(s+2)}^{J}\right.$
@@
    In the smeared Ds–D(s+4) intersections, one often encounters two different types of the Ramond–Ramond fields sources by different types of branes. This corresponds to $k=k^{\prime}=1$ in (3.3). Examples include D2–D6, D4–D8, and D3–D7 systems [51]. In all these cases, the fields $\psi$ and ${\bar{\psi}}$ are not independent since they come from a single dilaton in string theory. We will discuss implications of this constraint once the solutions are constructed below.
@@
    The D1–D5 is usually described in terms of a single three–form field strength that has both electric and magnetic components. Since in this article we are focusing on electric solutions, we will use an alternative description in terms of $F_{(3)}$ and $F_{(7)}$ with $k=k^{\prime}=1$ in (3.3). As for other Ds–D(s+4) systems, the scalars are constrained. The same action, although with different coupling to scalars, describes fundamental strings dissolved in NS5 branes.
@@
    The supergravity action for heterotic strings contains a three–form $H_{3}$ and 16 abelian vector fields. This system is covered by (3.3) with $(k,k^{\prime})=(16,1)$. In this case, ${\bar{\gamma}}=0$ and scalars ${\bar{\psi}}_{I}$ are turned off. Reduction of the action (3.3) to lower dimensions yields a system of vector fields coupled to scalars, which has been extensively studied in the past [7].
@@
Solution (3.1) describes a single stack of $p$–branes, so it is symmetric under $SO(n+1)$ rotations with $n=d-p-2$. Similarly, ignoring the matter fields in the first line of (3.3), one can construct a geometry describing a single stack of $s$–branes, which is invariant under $SO(n^{\prime}+1)$ rotations with $n^{\prime}=d-s-2$. On the other hand, one can uniformly smear the $s$–branes over a $(p-s)$–dimensional hyperplane. This would produce a geometry $SO(n+1)$ invariance, but with $(p+1)$ flat directions. Notice that since some of these directions are transverse to the $s$–branes, and other are longitudinal, there would be no rotational invariance in the $p$ spacial directions which was present in (3.1). Once the smeared branes with charges from the second line of (3.3) are combined with the $p$–branes charged under the fields in the first line, one arrives at a generalization of the ansatz (3.1):
@@
Note that in (3.3) we didn’t assume specific forms of the gauge potentials, but rather started with the most general field strengths consistent with symmetries and Bianchi identities. Once the Einstein’s equations fixed various powers in the metric and scalar fields in (3.3), integration of the Bianchi identities yielded very simple expressions for $(A_{(p+1)}^{I},{\bar{A}}_{(s+1)}^{J})$. The charges $(q_{I},{\bar{q}}_{J})$ contain an additional factor of $(n-1)$ in comparison to $(q^{\prime}_{I},{\bar{q}}^{\prime}_{J})$. We chose to write (3.3) in terms of two free parameters $(\nu,,{\bar{\nu}})$, but they are fixed in terms of the coefficients $(w_{1},w_{2},{\bar{w}}_{1},{\bar{w}}_{2},w_{3})$ in the action (3.3). In fact, to accommodate solutions of the form (3.3), these coefficients must be constrained: in contrast to the single stack of $p$–branes discusses in section 3.1, where two parameters $(w_{1},w_{2})$ in the action could have been traded for $\nu$ and rescaling of the scalar field, now such rescalings and changes in $(\nu,{\bar{\nu}})$ can absorb only four of the five constants. The parameters in the action are given by
@@
With these conditions, the near–horizon geometry of (3.3) reduces to $AdS_{s+2}\times S^{n}\times T^{p-s}$, and as expected, all scalars $(\psi_{I},{\bar{\psi}}_{J})$ saturate to constant values. As in section 3.2, one can go to the global coordinates on $AdS_{s+2}$ but unfortunately the procedure for a potential a potential term for the scalars and constructing geometries interpolating between $AdS_{s+2}\times S^{n}\times T^{p-s}$ and AdSd at infinity fails due to the presence of a torus. In the next section we will use an alternative method for constructing asymptotic AdSd with tori, but our example will use truncated versions of the action (3.3).
@@ 3.4 The most general solution for static $p$–branes
So far in this article we have considered various actions for gravity coupled to scalars and fluxes, and then imposed certain ansatze for the geometries. It is interesting to whether one can find more general solutions of equations of motion, and in this section we will investigate such possibilities. To avoid unnecessary clutter, we will focus on the action (3.1) with one vector field and one scalar, but qualitative conclusions will hold in more general cases as well. While the solution (3.1) for $k=1$ has only one charge parameter $q_{1}$, which appears in both the scalar and the field strength, we will show that the most general solution depends on two charges. A possibility of such extension was mentioned already in the original article by Horowitz and Strominger [1], and explicit solutions for two–charge black holes were constructed in [52, 53]. We will extend this discussion to branes and demonstrate that two–charged geometries give the most general static solutions.
In this subsection, we will focus on a simplified version of the action (3.1) [^12] Normalization of the kinetic term for $\psi$ is inspired by the normalization of the dilaton in string theory.:
$S=\int dx\sqrt{-g}\left[R-\frac{4}{d-2}(\partial\psi)^{2}-\frac{e^{\gamma\psi}}{2}F_{(p+2)}\wedge\star F_{(p+2)}\right]\,.$ (3.50)
@@
    The parameter $w$ defined by (C.32). Note that the relation between $C$ and $R$ is exact, and it is a consequence of the gauge choice which has already been made in (1). Expansions similar to (2) can be obtained by all values of $(p,n)$, and the absence of $q$ dependence in expressions for $(A,R,C,\phi)$ in terms of $B$ persists to all orders in $q$. Furthermore, the charge appears in the expression for $B$ only through an expansion of a simple logarithm.
@@
Solution (3)–(3) describes a geometry of $p$–branes that depends on four parameters, $(a_{1},a_{2},c,q)$. A closely related two–parameter solution describing black holes in four dimensions was recently constructed in [53], and now we will recover these four–dimensional geometries as special cases of our solution.

### v2

@@ 3.1 Solution for a single stack of multi–charged $p$–branes
$S=\int d^{d}x\sqrt{-g}\left[R+\frac{N_{p}\gamma^{2}\nu}{4[n_{-}\nu-2]}(\partial\psi)^{2}-\frac{e^{\gamma\psi}}{2(p+2)!}(F_{(p+2)})^{2}\right]\,.$ (3.7)
@@
    Assuming that $n_{-}\nu\neq 2$, one can rescale the scalar field to choose a convenient normalization. There are two natural options: the canonical normalization that would give
@@
    We will choose the latter option, which gives the kinetic term $\frac{4}{d-2}(\partial\psi)^{2}$. Then comparing (3.7) with the action of type IIB supergravity [106, 107, 108, 109], one can embed fundamental strings and various branes:
@@
    Since the geometry (3.1) has multiple gauge fields, it covers not only individual branes, but some smeared brane intersections as well. Specifically, let us consider a configuration D(p+4) branes stretched in the directions $(t,x^{1},\dots,x^{p+4})$ and Dp branes stretched in $(t,x^{1},\dots,x^{p})$. In the string frame, the metric reads
@@
    Upon smearing the lower dimensional branes in the $(x^{p+1},\dots,x^{p+4})$ directions and reducing the system along the four–dimensional torus spanned by these coordinates, one ends up with an effective supergravity action (3.1) with two gauge fields: one is sourced by D(p+4) branes, and another one is sourced by the Dp branes. Generically, there are also two scalars: the original dilaton and the volume factor of the 4–torus. Dimensional reduction from type II supergravity gives
@@
    $F_{5}$ $=$ $\frac{1}{4}dt\wedge d\left[\frac{1}{h_{1}}dX_{123}+\frac{1}{h_{2}}dX_{1}dY_{23}-\frac{1}{h_{3}}dX_{2}dY_{13}+\frac{1}{h_{4}}dX_{3}dY_{12}\right]+dual,$
    $h_{i}$ $=$ $1+\frac{Q_{i}}{r},\quad f=h_{1}h_{2}h_{3}h_{4}\,.$ (3.21)
@@
    we conclude that the scalars are subject to the constraint $\sum\psi_{i}=0$ due to the relation $k\nu n_{-}=2$. As we will show in the next subsection, this has important implications for the near–horizon limit of the geometry (f).
@@ 3.2 Branes in the near–horizon limit and wormhole geometries
In the last subsection we constructed the geometries (3.1) which describe multi–charged black branes with regular horizons located at $r_{h}=2m$. As in the Schwarzschild case, one encounters a coordinate singularity at $r=r_{h}$, and the space can be continued using Kruskal coordinates. In this subsection we will focus on the extremal case of the geometries (3.1), for which the horizon is pushed to $r=0$, and explore their near–horizon limits. We will see that for a particular choice of the parameter $\nu$, such near horizon limit gives a product of the $n$–dimensional sphere and an AdS space on the Poincare patch according to the same mechanism that led to the discovery of the AdS/CFT correspondence [12]. Going to the global coordinates on AdS, one removes the coordinate singularity at the origin, but also cuts out the asymptotically–flat region [13, 14, 15]. However, such a transition also allows one to introduce a cosmological constant or a nontrivial potential for scalars and extend the solutions (3.1) to geometries with AdSp+n+2 rather than flat asymptotics. In this subsection we will explore such modifications, generalizing the singly-charged wormhole–type geometries which were constructed in [42].
@@
The fluxes are still given by (3.1). The solutions (3.2) have curvature singularities at $r=0$ unless $k\nu n_{-}=2$: the easiest way to see this is to observe that the scalars make divergent contributions to the stress–energy tensor, but singularities can also be found by direct calculations. The regular solutions with $k\nu n_{-}=2$ have the following properties:
@@
    which holds everywhere, not only in the near horizon limit. Therefore, we conclude that to avoid curvature singularities in the solutions, the action (3.1) must have $k$ gauge fields but only $k-1$ independent scalars.
@@
    Some string theory examples of geometries with the regularity condition $k\nu n_{-}=2$ were reviewed in section 3.1, and not surprisingly, they contain the extremal D3, M2, M5 branes, D1–D5 system, and D3–D3–D3–D3 intersections. It would be interesting to find physical applications of other regular cases as well. Note that many string theory solutions that do not satisfy the regularity condition $k\nu n_{-}=2$, such as geometries produced by fundamental strings and D branes, are allowed in the full theory and lead to singularities only in the supergravity approximation.
@@
We begin with reviewing the results for the $k=1$ case, which were obtained in [42]. In this case, there is only one gauge field and no scalars due to the constant (3.27) so the potential is representing the genuine cosmological constant everywhere. Both the geometry (3.29) and the asymptotic AdSp+n+2 have the isometries of an $n$–dimensional sphere and a $p+1$–dimensional AdS space, so we will assume that the connecting solution has these isometries as well. This leads to an ansatz for the metric and the field strength:
@@
As demonstrated above, regular geometries can be obtained only if the condition $k\nu n_{-}=2$ is satisfied, and this imposes the constraint (3.27) on scalar fields. For such configurations, the ansatz (3.31) has a unique extension consistent with symmetries:
@@
Index $I$ runs from $1$ to $(k-1)$, and the values of $e^{(I)}_{0}$ cab be extracted from (3.26). These $(k+1)$ expansions can be substituted into three Einstein’s equations and $(k-1)$ equations for the scalars, and as in the case of one charge (3.31), one of Einstein’s equations turns out to be redundant. Therefore, at each order in $r$, one finds $(k+1)$ linear equations for $(k+1)$ variables, leading to the unique solution. The only exception from this rule is zeroth order, where one encounters a nonlinear constraint on $(A_{0},C_{0})$, which is solved by a straightforward counterpart of (3.33):
@@
In particular, one concludes that solutions (3.35) regular at $r=0$ are uniquely specified by one parameter $C_{0}$. Beyond the perturbative expansions (3.36) one can perform numerical integration of the system and analyze the behavior of solutions at infinity. One finds a phase structure which is qualitatively similar to items (c)–(f) above with one major difference: now the scalar fields $\psi_{I}$ have nontrivial profiles. We conclude that an asymptotic behavior from each of the options (d)-(f) above leads to a unique regular solution (3.35). The numerical values $(C_{\star},C_{\bullet})$ depend on the number of scalar fields and values of charges.
@@ 3.3 Smeared branes and brane intersections
In section 3.1 we constructed geometries (3.1) which solve equations of motion coming from the action (3.1). To do so, we imposed the ansatz (3.1) in which the number of flat directions was correlated with the rank of the field strength. Specifically, all flat directions were longitudinal to the branes. In this subsection we will extend the earlier results in two directions: we will construct the geometries describing smeared branes, in which some of the flat directions are transverse to the worldvolume, and the solutions produced by branes within branes. Some examples of the second type of solutions were briefly discussed in section 3.1, where we analyzed their reductions to lower dimensions, and here we will construct the full uplifted geometries. In contrast to the string theory examples discussed in section 3.1, generically we will be able to embed $s$–branes inside $p$ branes for any values of $(s,p)$, and the well known relation $p=s+4$ is recovered when the couplings of scalars are adjusted to their string theoretic values.
@@
$S$ $=$ $\int d^{d}x\sqrt{-g}\left[R-\sum_{I}^{k}\frac{e^{\gamma\psi_{I}}}{2(p+2)!}(F_{(p+2)}^{I})^{2}-\sum_{J}^{\bar{k}}\frac{e^{\bar{\gamma}{\bar{\psi}}_{J}}}{2(s+2)!}({\bar{F}}_{(s+2)}^{J})^{2}\right.$
@@
    In the smeared Ds–D(s+4) intersections, one often encounters two different types of the Ramond–Ramond fields sourced by different types of branes. This corresponds to $k=k^{\prime}=1$ in (3.3). Examples include D2–D6, D4–D8, and D3–D7 systems [115, 116, 117, 118, 119]. In all these cases, the fields $\psi$ and ${\bar{\psi}}$ are not independent since they come from a single dilaton in string theory. We will discuss implications of this constraint once the solutions are constructed below.
@@
    The D1–D5 system is usually described in terms of a single three–form field strength that has both electric and magnetic components. Since in this article we are focusing on electric solutions, we will use an alternative description in terms of $F_{(3)}$ and $F_{(7)}$ with $k=k^{\prime}=1$ in (3.3). As for other Ds–D(s+4) systems, the scalars are constrained. The same action, although with different coupling to scalars, describes fundamental strings dissolved in NS5 branes.
@@
    The supergravity action for heterotic strings contains a three–form $H_{3}$ and 16 abelian vector fields. This system is covered by (3.3) with $(k,{\bar{k}})=(1,16)$. In this case, ${\bar{\gamma}}=0$ and scalars ${\bar{\psi}}_{I}$ are turned off. Reduction of the action (3.3) to lower dimensions yields a system of vector fields coupled to scalars, which has been extensively studied in the past [16, 17, 18, 19, 20, 21].
@@
Solution (3.1) describes a single stack of $p$–branes, so it is symmetric under $SO(n+1)$ rotations with $n=d-p-2$. Similarly, one can construct a geometry describing a single stack of $s$–branes, which is invariant under $SO(n^{\prime}+1)$ rotations with $n^{\prime}=d-s-2$. On the other hand, one can uniformly smear the $s$–branes over a $(p-s)$–dimensional hyperplane. This would produce a geometry with $SO(n+1)$ invariance and $(p+1)$ flat directions. Notice that since some of these directions are transverse to the $s$–branes, and others are longitudinal, there would be no rotational invariance in the $p$ spacial directions which was present in (3.1). Once the smeared branes with charges from the second line of (3.3) are combined with the $p$–branes charged under the fields in the first line, one arrives at a generalization of the ansatz (3.1):
@@
Note that in (3.3) we did not assume specific forms of the gauge potentials, but rather started with the most general field strengths consistent with symmetries and Bianchi identities. Once the Einstein’s equations fixed various powers in the metric and scalar fields in (3.3), integration of the Bianchi identities yielded very simple expressions for $(A_{(p+1)}^{I},{\bar{A}}_{(s+1)}^{J})$. The charges $(q_{I},{\bar{q}}_{J})$ contain an additional factor of $(n-1)$ in comparison to $(q^{\prime}_{I},{\bar{q}}^{\prime}_{J})$. We chose to write (3.3) in terms of two free parameters $(\nu,,{\bar{\nu}})$, but they are fixed in terms of the coefficients $(w_{1},w_{2},{\bar{w}}_{1},{\bar{w}}_{2},w_{3})$ in the action (3.3). In fact, to accommodate solutions of the form (3.3), these coefficients must be constrained: in contrast to the single stack of $p$–branes discusses in section 3.1, where two parameters $(w_{1},w_{2})$ in the action could have been traded for $\nu$ and rescaling of the scalar field, now such rescalings and changes in $(\nu,{\bar{\nu}})$ can absorb only four of the five constants. The parameters in the action are given by
@@
With these conditions, the near–horizon geometry of (3.3) reduces to $AdS_{s+2}\times S^{n}\times T^{p-s}$, and as expected, all scalars $(\psi_{I},{\bar{\psi}}_{J})$ saturate to constant values. As in section 3.2, one can go to the global coordinates on $AdS_{s+2}$ but unfortunately the procedure for finding the potential term for the scalars and constructing geometries interpolating between $AdS_{s+2}\times S^{n}\times T^{p-s}$ and AdSd at infinity fails due to the presence of a torus. In the next section we will use an alternative method for constructing asymptotic AdSd with tori, but our example will use truncated versions of the action (3.3).
@@ 3.4 The most general solution for static $p$–branes
So far in this article we have considered various actions for gravity coupled to scalars and fluxes, and then imposed certain ansatze for the geometries. It is interesting to see whether one can find more general solutions of equations of motion, and in this section we will investigate such possibilities. To avoid unnecessary clutter, we will focus on the action (3.1) with one vector field and one scalar, but qualitative conclusions will hold in more general cases as well. While the solution (3.1) for $k=1$ has only one charge parameter $q_{1}$, which appears in both the scalar and the field strength, we will show that the most general solution depends on two charges. A possibility of such extension was mentioned already in the original article by Horowitz and Strominger [1], and explicit solutions for two–charge black holes were constructed in [120, 121]. We will extend this discussion to branes and demonstrate that two–charged geometries give the most general static solutions.
In this subsection we will focus on a simplified version of the action (3.1) [^12] Normalization of the kinetic term for $\psi$ is inspired by the normalization of the dilaton in string theory.:
$S=\int d^{d}x\sqrt{-g}\left[R-\frac{4}{d-2}(\partial\psi)^{2}-\frac{e^{\gamma\psi}}{2(p+2)!}(F_{(p+2)})^{2}\right]\,.$ (3.50)
@@
    The parameter $w$ is defined by (C.32). Note that the relation between $C$ and $R$ is exact, and it is a consequence of the gauge choice which has already been made in (1). Expansions similar to (2) can be obtained for all values of $(p,n)$, and the absence of $q$ dependence in the expressions for $(A,R,C,\phi)$ in terms of $B$ persists to all orders in $q$. Furthermore, the charge appears in the expression for $B$ only through an expansion of a simple logarithm.
@@
Solution (3)–(3) describes a geometry of $p$–branes that depends on four parameters, $(a_{1},a_{2},c,q)$. A closely related solution describing black holes in four dimensions was recently constructed in [121], and now we will recover these four–dimensional geometries as special cases of our solution.